Corporate Tax Rates as of 2015

(Table 2.3 of textbook)

Taxable Income

Pay this amount on base income

Plus this percentage on anything over the base

$0 – $50,000

$0

15%

$50,001 – $75,000

$7,500

25%

$75,001 – $100,000

$13,750

34%

$100,001 – $335,000

$22,250

39%

$335,001 – $10,000,000

$113,900

34%

$10,000,001 – $15,000,000

$3,400,000

35%

$15,000,000 – $18,333,333

$5,150,000

38%

Over $18,333,333

$6,416,667

35%

Greatland, Inc. had $284,000 in 2015 taxable income. Using the tax schedule from Table 2.3, what is the company's 2015 income taxes, average tax rate, and marginal tax rate, respectively? A. $110,760, 39.0%, 39.0%

B. $96,560, 34.0%, 34.0%

C. $94,010, 34.0%, 39.0%

D. $94,010, 39.0%, 33.1%

E. $94,010, 33.1%, 39.0%

15. You are considering a stock investment in one of two firms (AllDebt, Inc. and AllEquity, Inc.), both of which operate in the same industry and have identical operating income of $600,000. AllDebt, Inc. finances its $1.2 million in assets with $1.0 million in debt (on which it pays 10 percent interest annually) and $0.2 million in equity. AllEquity, Inc. finances its $1.2 million in assets with no debt and $1.2 million in equity. Both firms pay a tax rate of 30 percent on their taxable income. What are the asset funders' (the debt holders and stockholders') resulting return on assets for the two firms? A. 37.5%, and 35.0%, respectively B. 37.5%, and 37.5%, respectively

C. 39.2%, and 35.0%, respectively

D. 37.5%, and 50.0%, respectively

E. 50.0%, and 50.0%, respectively
Answer #1

1,

Total Taxable income = $284000

This lies in the slab 100001 - 335000

Hence, tax = 22250 + 39% of (284000 - 100000) = 22250 + 39% of 184000 = 22250 + 71760 = $94010

Average tax rate = total tax / Total income *100% = 94010 / 284000 * 100% = 33.10%

Marginal Tax rate = tax rate of slab = 39%

Hence, correct option is E. $94,010, 33.1%, 39.0%